OpenStudy (lillian_a):

5. A rancher is building a rectangular enclosure. He has to build only 3 sides of the enclosure as a river forms the 4th side. If he has 600 feet of fence, what’s the largest area he can enclose?

OpenStudy (anonymous):

Let w and L be the width and length respectively. L = 600 - 2w Area = w L = w(600 - 2w) Set the derivative of w(600 - 2w) to zero and solve for w. Then calculate L using the new w.